• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Ole Hansen

    Log entry time 00:16:13 on February27,2009

    Entry number 263602

    keyword=Swing Shift Summary

    Shift Workers:

    Ole Hansen (SL)
    Yunxiu Ye (TO)
    Dustin McNulty (3rd)

    Summary:
    Beam was restored after pass change to 5-pass shortly after the start of the shift. We performed the beam checkout procedure. Noticed that the maximum raster amplitude in y at this energy is 3mm, so we continued with a raster size of 3.5x3.0 mm^2. Took reference cell data with N2 for kinematics M, then switched to kinematics K, took reference cell data, then moved on to production with 3He for these kinematics. Beam disappeaed for 45 minutes starting around 21:50 due to rf problems. Beam eventually came back, but availability was still not good for the remainder of the shift (trips for 5-10 minutes at a time).

    Runs:

    LHRS:

    Run #events Comments
    20425-20429 -- Junk, checkout, spot checks
    20430 112k kine M, N2 reference cell run
    20431 142k kine K, N2 reference cell run
    20432 O(1e5) kine K, 3He production
    20433 in progress

    BigBite:

    1840-1841 -- junk, checkout
    1842-1843 -- prescale factor tests
    1844 7M ref cell production
    1845 7M
    1846 ?
    1847 320k 3He production
    1848 7M
    1849 6M end run because of long beam down
    1850 7M
    1851 in progress